OpenStudy (anonymous):

integral of exp(-kt) from 0 to infinity

OpenStudy (xapproachesinfinity):

That indefinite integral, Use the property (exp(t)) '=exp() t

OpenStudy (xapproachesinfinity):

Sorry I'm using the phone, things not going well lol

OpenStudy (anonymous):

Thank you, but I still don't get it?

OpenStudy (xapproachesinfinity):

Actually you have to to use this property (expkt)'=kexp(t)

OpenStudy (anonymous):

Thanks, I will try that.

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty }e ^{-kt }dt=\frac{ e ^{-kt} }{ -k }|~from~0~\to~\infty =\frac{ -1 }{ k }\frac{ 1 }{ e ^{k t} } |~0~\to~\infty \] \[=\frac{ -1 }{ k }\left( 0-\frac{ 1 }{ e ^{0} } \right)=\frac{ 1 }{ k }\]

OpenStudy (anonymous):

Thank you very much, that is straightforward and clear.

OpenStudy (anonymous):

yw